I have no idea how to do this. Could someone work this out for me.
2x-3y+4z=10
5y-2z=-16
3z=9

Answer by Alan3354(69443)

2x-3y+4z=10
5y-2z=-16
3z=9
-----------
Since the last eqn makes it easy, use substition
2z = 9
z = 3
------
2x -5y-2z = -16
5y-6 = -16
5y = -10
y = -2
--------
2x-3y+4z = 10
2x + 6 + 12 = 10
2x = -8
x = -4
------------
Most systems of 3 are more tedious.